If P(A) = 0.45, P(B) = 0.35, and P(A and B) = 0.25, then P(A | B) is
1
Expert's answer
2014-09-19T09:05:15-0400
P(A | B)=P(A and B)/P(B)=0.25/0.35=25/35=5/7.